# SimCity: Multi-Agent Urban Development Simulation with Rich Interactions

### v1

@@ 5.1 Emergence of Macroeconomic Phenomena (RQ1)
The Phillips Curve (Phelps, 1967) describes the short-run inverse relationship between unemployment and inflation, while Okun’s Law (Okun, 1963) depicts the quarterly negative relationship between changes in the unemployment rate and real GDP growth. Following prior work (Li et al., 2024), we use these well-known empirical regularities as major tests of the plausibility of macroeconomic dynamics in our simulation. Recent economic studies using more modern methodologies and granular data document that the Phillips Curve slope is small and has flattened in recent decades (Hazell et al., 2022; Furlanetto & Lepetit, 2024). Consequently, a steeper or more negative slope does not necessarily indicate a more faithful reproduction of the Phillips Curve. We plot the relationships using one data point per quarter in Figure 2131313For simulations that operate on an annual time scale, we use annual data points instead.. As shown in Figure 2, SimCity successfully demonstrates these relationships. [^14] We omit AI-Economist from this comparison because unemployment is not incorporated in its framework.
Figure 2: Emergence of the Phillips Curve and Okun’s Law in SimCity simulations. $r$-value is the Pearson correlation coefficient and $p$-value indicates the statistical significance of it.
Figure 3: Beveridge Curve and other macroeconomic emergences from SimCity.
@@
The Beveridge Curve captures the negative relationship between the job vacancy rate, which is the number of unfilled job openings as a proportion of total job postings, and the unemployment rate (Blanchard & Diamond, 1989; Lengnick, 2013). SimCity successfully exhibits this relationship too, whereas other frameworks in comparison lack the firm module needed to explicitly model job vacancies.
Further Tests
We further validate our model by examining Engel’s Law, a foundational empirical regularity, and the Law of Demand, a fundamental economic principle. Engel’s Law states that the proportion of income spent on food (the Engel coefficient) decreases as income rises (Chai & Moneta, 2010). Simulations from SimCity correctly exhibit this pattern, reflecting that LLM-driven agents display human-like preferences. Notably, validating Engel’s Law requires a heterogeneous goods market, which previous frameworks lacked and therefore could not support.
To test the Law of Demand, we estimate the Price Elasticity of Demand (PED) for each good, defined as $E=\frac{\Delta\log Q}{\Delta\log P}$, which measures how quantity demand responds to changes in price. Standard economic theory predicts a negative elasticity for normal goods, but the magnitude depends on the importance of the good to consumers (Mas-Colell et al., 1995). As shown in Figure 3, SimCity not only exhibits the general negative association predicted by the Law of Demand but also captures the variation in elasticity across goods. For example, necessities like “food/beverages” are more inelastic (1<E<0-1<E<0), whereas goods such as “motor/vehicles/trailers/semitrailers” are more elastic (E<1E<-1), which is a pattern confirmed by numerous empirical studies (Perloff, 2009; Nelson et al., 2014). Last but not least, as shown in Figure 3 (top-right), our volatility of investment is larger than that of consumption, which matches common-sense.
5.2 Robustness (RQ2)
@@
5.3 Urban Expansion During the Move-In Phase (RQ3)
As mentioned in Section 3.1, our simulation involves a phase in which households move in. Figure 5 illustrates the developments during this stage. As shown in the figure, in the first 15 months, GDP steadily increases with the influx of residents. Then, as the influx ceases, the city transitions from a period of expansion to one of stable development after more than ten months of fluctuations. Meanwhile, we note that the VLM, without additional prompts, forms a clustered structure with residential areas near the center and production buildings on the periphery.
5.4 Case Study: Exogenous Price Impulse (RQ4)
We separately apply a price-down and a price-up impact on the city. In the beginning of year 15, we randomly select 7 of 44 goods in two different simulations and apply a 50% price-down and a price-up impulse separately. We then let the simulation run for 6 years. Price-down impulse response is shown as Figure 6 and price-up at Figure 7 (Appendix B.2). The experiment shows that the good’s price level gradually returns to equilibrium after an exogenous shock. Due to the effect of price stickiness, the response to the exogenous shock will take some time. Note that dotted lines in the figure represent long-run average prices of goods. Since the shock is one time and the fundamentals of the economy did not change, prices revert close to their original levels in the long-run.
@@ 6 Limitations
Our agent’s decision-making is powered by a LLM. A significant challenge arises from the inherent difficulty in imposing robust constraints on the LLMs’ decisions. Furthermore, there exists a non-trivial probability that the LLM may generate highly aberrant or unexpected decisions, such as altering a price by several orders of magnitude. While the likelihood of any single such event is small, its probability becomes non-negligible over hundreds of iterations. In this work we apply simple heuristic-based checks prevent those egregious behavior.
@@ 7 Conclusion
With SimCity, we conduct complex simulations that incorporates four distinct LLM-driven economic roles (households, firms, the central bank and the government), with a visualized map and markets with heterogeneous goods. Experiments show that SimCity robustly matches a checklist of established macroeconomic phenomena and exhibit plausible responses to exogenous shocks. We hope SimCity could serve as a solid base platform towards more a realistic economic simulation.
@@ Acknowledgement
We sincerely thank Yixin Tao, Jian Li and Kaifeng Lyu for their valuable discussions and insightful comments that help improve this work.
@@ 8 Ethics Statement
This work utilize LLM-driven agents to simulate. No sensitive or private data is used. Names of agents are generated by LLMs and are not intended to represent any real individuals or entities. Our simulation has not been quantitatively calibrated with real-world data and does not constitute a prediction of reality.
@@ 9 Reproducibility Statement
To help with the reproducibility of our work, we provide all LLMs prompts in Appendix E.1. We call LLMs API with default sampling parameters. We repeat experiments with different random seeds (Appendix B.1) to ensure that our main results are robust. We plan to release code in near future.
@@ Appendix A Statement on LLM Usage
This work utilize LLM-driven agents to simulate. In addition, we use LLMs to improve the grammar and readability of this manuscript. The LLM was not used for any other scientific aspects of this work, and all intellectual content is solely the product of the authors.
@@ Appendix B Experiment Detail And Additional Results
Our simulation costs about 800,000 tokens, which is roughly \$0.25 per step. For a standard 180 steps simulation, the total cost is around \$180.
@@ B.2 Price Impulse
The price-up impulse response, Figure 7, is shown here due to space limitation.
@@ C.1 Notations
  • The goods set, $\mathcal{G}$, where $|\mathcal{G}|$ denotes the number of heterogeneous goods present in our simulation. Certain subsets of $\mathcal{G}$ have a specific meaning, such as $\mathcal{G}_{E}$ for essential goods, $\mathcal{G}_{T}$ for transportation goods and $\mathcal{G}_{D}$ for durable capital.
@@ C.2 Households
  • Labor Participation: An indicator $l_{i}\in\{0,1\}$, where 1 means the household is employed to work in a position denoted by $\mathrm{position}_{j}$ and receives a salary $S_{j}$, and 0 means the household is unemployed. Note we are modeling the extensive margin of labor, instead of the intensive margin (hours worked).
@@ C.3 Firms
    $D_{j}=\frac{\textbf{share}_{i,j}\,d_{i}\,m_{i}}{\sum\textbf{share}_{i}}.$
@@
    $\mathcal{V}_{i}=\frac{\sum_{j=t-12}^{t-1}\text{profit}_{j}}{12\cdot\max(\mathcal{I}_{d,t},\epsilon_{d})}+m_{i}-L_{i}+\mathcal{K}_{i}\cdot$
    where $\mathcal{I}_{d,t}$ denotes the GDP deflator, and $\epsilon_{d}=0.02$ is a lower bound introduced to prevent division by zero.
  • Loans: The amount of loan the firm owes to the financial system.
If a firm experiences zero income for a year and its assets are insufficient to cover its overdue debts, the firm is declared bankrupt. In that case, all employees are terminated, and the land it occupied is released.
@@ C.5 Central Bank and Financial System
Empirically, interest rate smoothing is a well-documented phenomenon. For example, Coibion & Gorodnichenko (2012) estimate a Taylor rule with smoothing and find that the smoothing coefficient typically ranges between 0.7 and 0.9 in developed economies. We choose $\rho=0.8$ as our parameterization.
@@ D.1 Environment Setup
There are 44 types of goods in SimCity. Each of them represents of an economic industry in the real world as categorized by Organization for Economic Co-operation and Development (OCED).
@@ D.2 Stages within Each Step
  • Metabolic Stage: New companies are established through equity financing from the investment pool, bankrupted companies are removed, and the population grows according to set rules.
@@ D.3 Render Module
We utilize flask-socketio to build a web-server and Vue.js to build a website.
@@ D.4 Profile Setup
To assign realistic initial cash holding to households, we estimate the income distribution using U.S. microdata. We use the 2023 American Community Survey (ACS) IPUMS microdata, which contains detailed household-level income information. We assume a lognormal distribution for household income, a standard approximation in the economics literature due to its simplicity and its ability to capture the right-skewed nature of income data. [^19] While some heavier-tailed distributions (e.g. Pareto) better approximate top incomes, our simulation does not focus on the super-rich. Hence, the lognormal distribution is a reasonable choice. See Hlasny (2021) for a recent summary of related work. Formally, the initial cash holding is drawn from
@@ D.5 Synthesis
Observing that the positions required skills in the generated firm templates contain many semantically similar skills, we use the prompt introduced in Listing 2 to call an LLM to merge these similar skills. Skills after merge is shown at Table 3.
@@ E.1 Prompting Structure
As mentioned in 4, the Agent will receive necessary information to make its decisions. The Profile includes some basic personal information about the Agent, such as current income, age, skills, job, and residence. The Report includes some of the Agent’s recent experiences and information. The Observation presents information about the overall environment that the Agent should be aware of, such as vacant positions, the unemployment rate, prices of goods and the Return on Investment (ROI).
@@ E.3 Full Examples
Your name is David, and you are 57 years old. Your ID is 2000111.You current have 1000.000 money in your account.
You are a Disorganized,Adventurous man
You are working at Construction Firm as Construction Worker with salary 1200.000. This work requires skills: Building, Physical Labor.
You are living in Residential Building at (5, 5) with rent 950. Your work place is at (3, 2). The distance between your home and your work place is 5.
You have experienced the following education:
- K-12
- Bachelor
You plan to spend a percentage of your income on the following items:
- Agriculture_Hunting_Forestry: 7.5
- Fishing_Aquaculture: 2.5
@@
- Motor_Vehicles_Trailers_SemiTrailers: 0.05
@@
Demand deposit interest rate is 37.9503%, and loan rate is 40.7381%.
Your loanable amount is 362000.0.
@@
This month, your income is 2500.000, which consists of the following parts:
- Salary: 1500.000(60.0000%)
- Welfare: 1000.000(40.0000%)
This month, your outcome is 2500.000, which consists of the following parts:
- Consumption: 1575.000(63.0000%)
- Food_Beverages: 900.000(57.1429%)
- Education: 150.000(9.5238%)
- Health_Social_Work: 450.000(28.5714%)
- Motor_Vehicles_Trailers_SemiTrailers: 75.000(4.7619%)
- Tax: 0.000(0.0000%)
- Rent: 925.000(37.0000%)
@@
- position_id: 2000130, Production Worker requires skills Operations Management, Physical Labor, with salary 300.000
@@
- position_id: 2000279, Production Worker requires skills Operations Management, Physical Labor, with salary 2000.000
You expect the work should cover your spending of 2500.000.
@@
- residential id: 200038 at (7, 3) with rent 975
@@
- residential id: 200010 at (11, 0) with rent 950
The social average return on investment is 11.7672%, and the interest rate of the bank is 37.9503%.
@@
- Food_Beverages: 81.344, 29.2090% (up)
- Textiles_Leather_Footwear: 75.000, 21.9512% (up)
- Wood_Cork_Products: 86.845, 22.3489% (up)
- Education: 65.748, 8.2185% (up)
- Health_Social_Work: 82.491, 19.0899% (up)
- Public_Administration_Defence_SocialSecurity: 14.000, -50.1743% (down)
- Motor_Vehicles_Trailers_SemiTrailers: 55.000, 325.8065% (up)
@@
During the past 9 months, your average income is 166.667:
- 9 months ago, you were unemployed, and your monthly income was 0.000.
- 8 months ago, you were unemployed, and your monthly income was 0.000.
@@
- 2 months ago, you were unemployed, and your monthly income was 0.000.
- 1 months ago, you were working at Construction Worker with salary 1500.000, and your monthly income was 1500.000.
@@
reasoning’: Currently, you are working as a Construction Worker earning a monthly income of 1500.000, which covers your basic living expenses but allows little room for savings or investments. Given that the average income in the past nine months was only about 166.667, your recent income increase is significant. However, with your current living expenses and attempts to pay for various goods, you are at a financial standstill. The available job positions include better-paying opportunities such as Production Worker with a salary of 2000.000, which will provide more financial relief. Therefore, I recommend you apply for a higher-paying position while reducing some expenditures to enhance your life situation.’,
@@
"modify_needs_percentage(’Food_Beverages’, 0.1)",
"modify_needs_percentage(’Health_Social_Work’, 0.1)"
@@
Loans:
- id: 369, loan amount: 13924261.356747806, should be paid back at Year 13 Month 1
Demand deposit interest rate is 56.0464%, and loan rate is 58.0693%.
Your loanable amount is 10000000.000000002.
@@
- Up to 10000 get 5.0% tax
- Up to 20000 get 10.0% tax
- Up to 50000 get 15.0% tax
- Up to 100000 get 20.0% tax
The government has collected 124774.65254701707 as tax.
@@
- 55701.10070718494 are from 100000 down to 0
- 0.0 are from down to 100000
@@
- 0.0 are from down to 100000
- 69073.55183983216 are from 100000 down to 0
The government has distributed 50000 to every citizen as UBI.
Therere 200 citizens in the city, and working or living in 122 buildings.
@@
- Consumption: 489137.5517
- Investment: 1299054.2543
and total nominal GDP is 1788191.8059999999.
The real GDP, calculated by the base years price, is 1810789.8210999998.
The GDP deflator is 0.9875203544681556.
The social equality defined as Gini coefficient of income is 0.8564, and the wealth inequality is 0.8576.
The unemployment rate is 0.235.
@@
- Year 12 Month 2: Consumption: 803850.0466, Investment: 1430793.5346, Government Expense: 50000.0, GDP: 2284643.5812, Wealth Gini Coefficient: 0.8619, Income Gini Coefficient: 0.7464, Unemployment Rate: 0.335, Homeless Rate: 0.095, Average Happiness: 0.9475, Average Wage: 1234.325, Average Position Salary: 1755.1287, Broad Money Supply: M0 22620930.0, M1 22620930.0, Deposit Rate: 0.4203, Loan Rate: 0.4421, Reserve Rate: 0.05
@@
strategies’: "The citys current balance indicates a significant amount of debt with no savings, suggesting that financial management needs immediate attention. The high unemployment rate points to economic distress, indicating a need for initiatives that stimulate job creation and investment. Adjusting tax brackets, especially for buildings, could also enhance government revenue, which is crucial given the existing loans. Additionally, reviewing the Universal Basic Income (UBI) distribution may free up resources for essential services.",
@@
"adjust_bracket(’Building’, [(10000, 0.04), (50000, 0.07), (100000, 0.12)])",
invest(10000000)’,
borrow(10000000)’,
adjust_UBI(40000)’
@@
- Nonenergy_Mining_Products: 47.915
- Food_Beverages: 81.344
@@
- Education: 65.748
- Health_Social_Work: 82.491
@@
Therere 200 citizens in the city, and 25 residential buildings have capacity, and provide 115 empty houses.
The homeless rate is 0.125, which means 25 citizens are homeless.
and the average rent is 958.6206896551724
@@
The size of investment pool is: 2103539.50867976
@@
- Electricity_Gas_Steam_AirConditioning: Lack 1051.8885135094476
- Mining_Support_Services: Lack 0.0

### v2

@@ 5.1 Macroeconomic Emergence (RQ1)
The Phillips Curve (Phelps, 1967) describes the short-run inverse relationship between unemployment and inflation, while Okun’s Law (Okun, 1963) depicts the quarterly negative relationship between changes in the unemployment rate and real GDP growth. Following prior work (Li et al., 2024), we use these well-known empirical regularities as major tests of the plausibility of macroeconomic dynamics in our simulation. Recent economic studies using more modern methodologies and granular data document that the Phillips Curve slope is small and has flattened in recent decades (Hazell et al., 2022; Furlanetto and Lepetit, 2024). Consequently, a steeper or more negative slope does not necessarily indicate a more faithful reproduction of the Phillips Curve. We plot the relationships using one data point per quarter in Figure 2333For simulations that operate on an annual time scale, we use annual data points instead.. As shown in Figure 2, SimCity successfully demonstrates these relationships. [^4] We omit AI-Economist from this comparison because unemployment is not incorporated in its framework.
@@
The Beveridge Curve captures the negative relationship between the job vacancy rate, which is the number of unfilled job openings as a proportion of total job postings, and the unemployment rate (Blanchard and Diamond, 1989; Lengnick, 2013). SimCity successfully exhibits this relationship, whereas other frameworks in comparison lack the firm module needed to explicitly model job vacancies.
Figure 4: Estimated Price Elasticity of Demand (PED) across different product categories in the SimCity environment, obeying the Law of Demand and commonsense notions of necessity.
Engel’s Law
Engel’s Law states that the proportion of income spent on food (the Engel coefficient) decreases as income rises (Chai and Moneta, 2010). Simulations from SimCity correctly exhibit this pattern, reflecting that LLM-driven agents display human-like preferences. Notably, validating Engel’s Law requires a heterogeneous goods market, which previous frameworks lacked and were therefore unable to study.
Law of Demand
To test the Law of Demand, we estimate the Price Elasticity of Demand (PED) for each good, defined as $E=\frac{\Delta\log Q}{\Delta\log P}$, which measures how quantity demand responds to changes in price. Standard economic theory predicts a negative elasticity for normal goods, but the magnitude depends on the importance of the good to consumers (Mas-Colell et al., 1995). As shown in Figure 4, SimCity not only exhibits the general negative association predicted by the Law of Demand but also captures the variation in elasticity across goods. For example, necessities like “food/beverages” are more inelastic (1<E<0-1<E<0), whereas goods such as “motor/vehicles/trailers/semitrailers” are more elastic (E<1E<-1), which is a pattern confirmed by numerous empirical studies (Perloff, 2009; Nelson et al., 2014). Lastly, as shown in Figure 3 (top-right), our volatility of investment is larger than that of consumption, which is consistent with real-world business cycle patterns.
5.2 Simulation Robustness (RQ2)
@@
5.3 Urban Expansion During Move-In (RQ3)
As mentioned in Section 3.1, our simulation involves a phase in which households move in. Figure 5 illustrates the developments during this stage. As shown in the figure, in the first 15 months, GDP steadily increases with the influx of residents. Then, as the influx ceases, the city transitions from a period of expansion to one of stable development after more than ten months of fluctuations. Meanwhile, we note that the VLM, without additional prompts, forms a clustered structure with residential areas near the center and production buildings on the periphery, in line with Alonso-Muth-Mills (AMM) model of the mono-centric city (Brueckner, 2011).
To evaluate the role of visual information in generating realistic urban forms, we conduct an ablation study to evaluate the necessity of the visual module by replacing the map images with text-based descriptions of the city layout. Experiments show that without visual grounding, the city evolves into unrealistic geometries, such as narrow linear formations, contrasting with the organic clusters formed by the VLM-empowered agents (visual comparison provided in Appendix B.3).
@@ 5.3 Urban Expansion During Move-In (RQ3)
5.4 Case Study: Exogenous Price Shock (RQ4)
To investigate whether our framework can be used to analyze the effects of sudden external shocks to the economy, we separately apply a price-down and a price-up impact on the city. At the beginning of year 15, we randomly select 7 of 44 goods in two different simulations and apply a 50% price-down and a price-up impulse separately. We then let the simulation run for 6 years. Conceptually, one can think of these sudden, ephemeral changes in price levels of the selected goods as coming from exogenous shocks in the world market. Take electricity prices as an example, which is one of the shock goods in our simulations, it could come from a significant but brief conflict among energy-producing countries.
Price-down impulse response is shown as Figure 6 and price-up at Figure 8 (Appendix B.2). The experiment shows that the good’s price level gradually returns to equilibrium after an exogenous shock. Due to price stickiness, the full effects of the exogenous shock take some time to develop. Importantly, note that dotted lines in the figure represent long-run average prices of goods. Since the shock is one-off and the fundamentals of the economy did not change, prices revert close to their original levels in the long-run, despite temporal dislocation. This stylized experiment showcases SimCity’s ability to produce qualitative predictions in novel economic scenarios.
@@ 6 Limitations
Our agent’s decision-making is powered by a LLM. A significant challenge arises from the inherent difficulty in imposing robust constraints on the LLMs’ decisions. Furthermore, there exists a non-trivial probability that the LLM may generate highly aberrant or unexpected decisions, such as altering a price by several orders of magnitude. While the likelihood of any single such event is small, its probability becomes non-negligible over hundreds of iterations. In this work, we apply simple heuristic-based checks to prevent such egregious behavior.
@@ 7 Conclusion
With SimCity, we conduct complex simulations that incorporate four distinct LLM-driven economic roles (households, firms, the central bank, and the government), with a visualized map and markets with heterogeneous goods. Experiments show that SimCity robustly matches a checklist of established macroeconomic phenomena and exhibits plausible responses to exogenous shocks. We hope SimCity could serve as a solid base platform for a more realistic economic simulation.
@@ Acknowledgement
We sincerely thank Jian Li and Kaifeng Lyu for their valuable discussions and insightful comments that help improve this work.
@@ 8 Ethics Statement
This work utilizes LLM-driven agents to simulate. No sensitive or private data is used. Names of agents are generated by LLMs and are not intended to represent any real individuals or entities. Our simulation has not been quantitatively calibrated with real-world data and does not constitute a prediction of reality.
@@ 9 Reproducibility Statement
To help with the reproducibility of our work, we provide all LLMs prompts in Appendix E.1. We call LLMs API with default sampling parameters. We repeat experiments with different random seeds (Appendix B.1) to ensure that our main results are robust. We plan to release code in the near future.
@@ Appendix A Statement on LLM Usage
This work utilizes LLM-driven agents to simulate. Additionally, we use LLMs to improve the grammar and readability of this manuscript. The LLM was not used for any other scientific aspects of this work, and all intellectual content is solely the product of the authors.
@@ Appendix B Experiment Detail And Additional Results
Our simulation costs about 800,000 tokens, which is roughly \$0.25 per step. For a standard 180-step simulation, the total cost is around \$180.
Base Year
We set prices for the first year as the base-year prices for calculating real GDP.
@@ B.2 Price Impulse
The price-up impulse response, Figure 8, is shown here due to space limitations.
@@
B.3 Ablation Study: Visual Module
To validate the contribution of the vision component, we compare the standard SimCity framework against a text-only variant. In this ablation setting, the establishment agent receives a list of building coordinates rather than a rendered map image to make location decisions.
As illustrated in Figure 9, the absence of the visual module significantly degrades the spatial logic of urban planning. While the VLM-based approach (Figure 9a) fosters a natural, concentric city structure with a distinct center, the text-only approach (Figure 9b) fails to interpret spatial density effectively. Consequently, the agents form an artificial linear arrangement of buildings (”two lines”), confirming that the visual module is crucial for interpreting complex spatial information that is difficult to convey through text alone.
(a) With VLM
(b) Text Description Only
Figure 9: Impact of the Visual Module on urban layout. The full model produces a realistic, clustered city center, whereas removing the visual input results in an unrealistic linear topology.
@@ C.1 Notations
  • The goods set, $\mathcal{G}$, where $|\mathcal{G}|$ denotes the number of heterogeneous goods present in our simulation. Certain subsets of $\mathcal{G}$ have a specific meaning, such as $\mathcal{G}_{E}$ for essential goods, $\mathcal{G}_{T}$ for transportation goods, and $\mathcal{G}_{D}$ for durable capital.
@@ C.2 Households
  • Labor Participation: An indicator $l_{i}\in\{0,1\}$, where 1 means the household is employed to work in a position denoted by $\mathrm{position}_{j}$ and receives a salary $S_{j}$, and 0 means the household is unemployed. Note that we are modeling the extensive margin of labor, rather than of the intensive margin (hours worked).
@@ C.3 Firms
    $D_{j}=\frac{\textbf{share}_{i,j}\,d_{i}\,m_{i}}{\sum_{k}\textbf{share}_{i,k}}.$
@@
    $\mathcal{V}_{i}=\frac{\sum_{j=t-12}^{t-1}\textbf{profit}_{j}}{12\cdot\max(\mathcal{I}_{d,t},\epsilon_{d})}+m_{i}-L_{i}+\mathcal{K}_{i}\cdot$
    where $\mathcal{I}_{d,t}$ denotes the GDP deflator, $\epsilon_{d}=0.02$ is a lower bound introduced to prevent division by zero and $\textbf{profit}_{j}$ denotes the net profit in month $j$, defined as the sales revenue minus the production expenditures.
  • Loans: The amount of loans the firm owes to the financial system.
If a firm experiences zero income for a year and its assets are insufficient to cover its overdue debts, the firm is declared bankrupt. In that case, all employees are terminated, and the land it occupies is released.
@@ C.5 Central Bank and Financial System
Empirically, interest rate smoothing is a well-documented phenomenon. For example, Coibion and Gorodnichenko (2012) estimates a Taylor rule with smoothing and finds that the smoothing coefficient typically ranges between 0.7 and 0.9 in developed economies. We choose $\rho=0.8$ as our parameterization.
@@ D.1 Environment Setup
There are 44 types of goods in SimCity. Each of them represents an economic industry in the real world as categorized by the Organization for Economic Co-operation and Development (OECD).
@@ D.2 Stages within Each Step
  • Metabolic Stage: New companies are established through equity financing from the investment pool, bankrupt companies are removed, and the population grows according to set rules.
@@ D.3 Render Module
We utilize flask-socketio to build a web server and Vue.js to build a website.
@@ D.4 Profile Setup
To assign realistic initial cash holdings to households, we estimate the income distribution using U.S. microdata. We use the 2023 American Community Survey (ACS) IPUMS microdata, which contains detailed household-level income information. We assume a lognormal distribution for household income, a standard approximation in the economics literature due to its simplicity and its ability to capture the right-skewed nature of income data. [^9] While some heavier-tailed distributions (e.g., Pareto) better approximate top incomes, our simulation does not focus on the super-rich. Hence, the lognormal distribution is a reasonable choice. See Hlasny (2021) for a recent summary of related work. Formally, the initial cash holding is drawn from
@@ D.5 Synthesis
Observing that the positions required skills in the generated firm templates contain many semantically similar skills, we use the prompt introduced in Listing 2 to call an LLM to merge these similar skills. Skills after merge are shown at Table 3.
@@ E.1 Prompting Structure
As mentioned in 4, the Agent will receive the necessary information to make its decisions. The Profile includes some basic personal information about the Agent, such as current income, age, skills, job, and residence. The Report includes some of the Agent’s recent experiences and information. The Observation presents information about the overall environment that the Agent should be aware of, such as vacant positions, the unemployment rate, prices of goods and the Return on Investment (ROI).
@@ E.3 Full Examples
Your name is David, and you are 54 years old. Your ID is 2000161.You current have 1878.741 money in your account.
You are a Disorganized,Unadventurous woman
You are currently unemployed.
You are living in Residential Building at (25, 15) with rent 450.
You plan to spend a percentage of your income on the following items, at a total of 1.11% of your income:
- Agriculture_Hunting_Forestry: 0.0
- Fishing_Aquaculture: 0.01
@@
- Motor_Vehicles_Trailers_SemiTrailers: 0.0
@@
Demand deposit interest rate is 0.6108%, and loan rate is 2.6108%.
Your loanable amount is 243757.48131444445.
@@
This month, your income is 2878.741, which consists of the following parts:
- Salary: 1000.000(34.7374%)
- Benefit: 1000.000(34.7374%)
- Profit: 0.000(0.0000%)
- Welfare: 878.741(30.5252%)
This month, your outcome is 1442.924, which consists of the following parts:
- Consumption: 1017.924(70.5459%)
- Agriculture_Hunting_Forestry: 297.363(29.2127%)
- Food_Beverages: 720.561(70.7873%)
- Rent: 425.000(29.4541%)
@@
- position_id: 2003486, Production Manager requires skills Management, Quality Control, with salary 600.000
@@
- position_id: 2000253, Plant Manager requires skills Management, Quality Control, with salary 5000.000
You expect the work should cover your spending of 1442.924.
@@
- residential id: 200089 at (4, 3) with rent 450
@@
- residential id: 200366 at (18, 19) with rent 450
The social average return on investment is 1.3949%, and the interest rate of the bank is 0.6108%.
@@
- Agriculture_Hunting_Forestry: 1325.850, 18.4137% (up)
- Fishing_Aquaculture: 187.039, 24.6316% (up)
- Food_Beverages: 3093.156, 20.1293% (up)
- Wood_Cork_Products: 995.833, 4.9505% (up)
- Accommodation_Food_Services: 309.991, 49.3516% (up)
- Education: 869.002, -10.6580% (down)
- Electricity_Gas_Steam_AirConditioning: 2046.818, 35.9158% (up)
- Telecommunications: 1277.031, 27.9131% (up)
@@
During the past 12 months, your average income is 444.167:
- 12 months ago, you were unemployed, and your monthly income was 0.000.
- 11 months ago, you were working at Customer Service Representative with salary 1400.000, and your monthly income was 0.000.
@@
- 1 months ago, you were working at Warehouse Worker with salary 1000.000, and your monthly income was 1000.000.
Here are the news in the past months:
1985-01-01 00:$00:$ A sudden collapse in food sector starts. All food related sectors production efficiency decreases to 50% of its original value.
@@
reasoning’: You are currently unemployed and have a low income relative to your expenses. Your current financial stability is precarious as your spending exceeds what you earn. The job market presents several opportunities, particularly in management roles, which match your skills. Since your rent is stable and there are affordable housing options available, finding a job with a salary that covers your expenses or investing in your skills can greatly improve your situation. Additionally, you should consider adjusting your spending on necessities to better reflect your current income.’,
@@
"modify_needs_percentage(’Food_Beverages’, 0.08)",
"modify_needs_percentage(’Accommodation_Food_Services’, 0.01)",
"modify_needs_percentage(’Health_Social_Work’, 0.05)"
@@
You have no loans.
Demand deposit interest rate is 0.6108%, and loan rate is 2.6108%.
Your loanable amount is 10000878.74065722.
Here are the news in the past months:
1985-01-01 00:$00:$ A sudden collapse in food sector starts. All food related sectors production efficiency decreases to 50% of its original value.
@@
- Down to 0.000 get 5.0% tax
- From 0.000 to 50000.000 get 10.0% tax
The government has collected 104553.065 as tax.
@@
- 0.000 are from down to 50000.000
- 33144.289 are from 50000.000 to 0.000
@@
- 0.000 are from down to 100000.000
- 71408.776 are from 100000.000 to 0.000
The government has distributed 64000.000 to every citizen as UBI.
Therere 400 citizens in the city, and working or living in 330 buildings.
@@
- Consumption: 1637604.621
- Investment: 332733.1813
- Government Expense: 0
and total nominal GDP is 1970337.8023.
The real GDP, calculated by last months price, is 1921755.1228.
The GDP deflator is 1.025280369451658.
The social equality defined as Gini coefficient of income is 0.8438, and the wealth inequality is 0.8353.
The unemployment rate is 0.3275.
@@
- Year 14, Month 10: Consumption: 1651127.9157, Investment: 274729.127, Government Expense: 0, GDP: 1925857.0427, Wealth Gini Coefficient: 0.8505, Income Gini Coefficient: 0.8086, Unemployment Rate: 0.285, Homeless Rate: 0.08, Average Happiness: 0.9555, Average Wage: 1732.675, Average Position Salary: 1722.0206, Broad Money Supply: M0 44111103.5414, M1 44111103.5414, Deposit Rate: 0.0428, Loan Rate: 0.0628, Reserve Rate: 0.05
- Year 14, Month 11: Consumption: 1596120.9542, Investment: 463245.2733, Government Expense: 0, GDP: 2059366.2275, Wealth Gini Coefficient: 0.8475, Income Gini Coefficient: 0.8213, Unemployment Rate: 0.2425, Homeless Rate: 0.1175, Average Happiness: 0.9522, Average Wage: 1736.1225, Average Position Salary: 1742.6331, Broad Money Supply: M0 44111103.5414, M1 44111103.5414, Deposit Rate: 0.0394, Loan Rate: 0.0594, Reserve Rate: 0.05
@@
strategies’: Given the recent collapse in the food sector, it is crucial to address food security and the consequent economic downturn. The high unemployment rate and income inequality suggest that social programs and job creation initiatives will be beneficial. Additionally, recognizing that the government has no savings and a balance suitable for moderate investments, we should focus on stabilizing the economy and enhancing living conditions.’,
@@
adjust_UBI(70000.00)’,
invest(500000.00)’,
"adjust_bracket(’Citizen’, [(0.000, 0.05), (50000.000, 0.07)])",
"adjust_bracket(’Building’, [(0.000, 0.05), (100000.000, 0.12)
@@
- Agriculture_Hunting_Forestry: 263.236
- Nonenergy_Mining_Products: 2.250
@@
- Arts_Entertainment_Recreation: 60.000
- Other_Service_Activities: 300.000
@@
Therere 1000 citizens in the city, and 125 residential buildings have capacity, and provide 532 empty houses.
The homeless rate is 0.132, which means 132 citizens are homeless.
and the average rent is 1403.5714285714287
@@
- Entertainment Production Studio(Produces Arts_Entertainment_Recreation): A studio that produces audiovisual content and provides entertainment recreation services. Cost 50000 to build, needs 7 employees to run
@@
The size of investment pool is: 5145622.093042261
@@
- Other_Transport_Equipment: Lack 0.0
- Electricity_Gas_Steam_AirConditioning: Lack 79.92203775782382
- Telecommunications: Lack 3914.2222825510344